3. (b) Solve the BVP z’y – 5xy + 8y = 24, y(0) = 3, y(1) = 0)
Answer #1

The Given tiff es is Xy_sky+ 8y = 24 test x=et - = logu X r = 018-1) where. I = 1 (25 - SKD+8) y = 24 [oco-1) -58 +8] +8] 4G=(2=-2 they Y = -x-12 + 3